100ml of 0.3 N HCl ismixed with 200ml of 0.6N H2SO4 what is the normality of H2S04 in final solution.

Grade:12

1 Answers

Arun
6 years ago
Dear student
 
 
Milli equivalents of HCl = 0.3*100 = 30 
 milli equivalents of H2SO4 = 200*0.6 = 120 
   
      30 milliequivalents of H2SO4 completely neutralizes 30 milliequivalents of HCl  
 
    normality =  milliequivalents/total volume 
                
                 N = 120-30/300
                  N = 0.3
 
 
Regards
Arun (askIITians forum expert)